# Adaptive Massively Parallel Coloring in Sparse Graphs

### v1

@@ 2 Sublinear LCA: High Level Challenges and Techniques
We design a deterministic LCA with a sublinear number of queries to find an acyclic orientation with a small outdegree $\beta=\Omega(\alpha)$ for a very large fraction of the nodes in a graph with arboricity $\alpha$. The informal statement is Lemma 1.1 and the formal one is Lemma 4.7.
@@ 2.1 Where should one query?
In order to steer the direction of the queries, consider the following coin dropping procedure. Give $n^{\delta}$ coins to node $v$, who naïvely (this approach will still not work) forwards the coins to all of its neighbors, giving each neighbor an equal share of the total number of $v$’s coins. Consider repeating this process recursively for every node holding coins, until coins cannot be further divided, which is in some sense a "volume-based" BFS. The forwarding of coins represents querying for nodes, and every node that has received a coin has been queried by $v$ and hence known to $v$. Optimally, all nodes of $D(v)$ receive a coin. Unfortunately, this approach fails for a large family of counter-examples: skewed dependency graphs (Figure 2b), comprising of a long path where every node has, e.g., $n^{\delta/10}$, neighbors in the lowest layer. In this case, $n^{\delta}$ coins will be already spent after only 10 steps of distributing coins, effectively learning only roughly $k\cdot n^{\delta/10}$ dependency graph nodes.
@@ 2.2 Our solution
Our solution uses the same coin dropping procedure, but we devise a more sophisticated forwarding rules that determine how to distribute the coins. These rules utilize the topology of the graph induced by $v$’s knowledge aka the nodes $v$ has discovered so far. The forwarding rules are also updated frequently during the course of the algorithm to steer the future exploration of the graph and making very strong usage of the adaptivity of the queries.
Recall that every node in $D(v)$ has at most $\beta$ neighbors outside of $D(v)$. Hence, when coins are at a node $u\in D(v)$ and $u$ forwards coins to $\beta+1$ neighbors, we ensure that at least one neighbor from $D(v)$ receives coins, or more precisely at least a $1/(\beta+1)$-fraction of the coins goes to neighbors in $D(v)$. If we keep recursively applying this approach to all nodes that still have coins available, this rough tells us, that after $k$ steps of forwarding, at least $x/(\beta+1)^{k}$ coins are at nodes that are contained in the dependency graph. Here $x$ is the number of coins that $v$ started with. However, it is not sufficient to choose $\beta+1$ neighbors arbitrarily, since it may be that we keep forwarding coins to nodes that we already know, i.e., nodes that have seen a coin before. In the end, it may be that all of these coins are just accumulated in a small part of $D(v)$ that we already know, or in other words, the process does not lead $v$ to querying new, undiscovered, parts of the dependency graph.
@@
Hence, we need to choose the correct neighbors to whom we forward the coins such that enough coins "flow" into the part of the dependency graph we have not yet learned, and not too many "leak out" outside of the dependency graph or "get stuck" in a part of the dependency graph that we already know. The rules for choosing the correct neighbors are chosen adaptively and frequently updated. Node $v$ keeps simulating an estimation of the $\beta$-partition (this estimation is called the induced $\beta$-partition, formalized in Definition 3.6) for the whole subgraph it has queried so far, which we call $S_{v}$ (which is only a small part of the whole graph). In the simulation, every node in $S_{v}$ is assigned a layer by $v$, or if no layer is assigned, we assume it is $\infty$. During an iteration, nodes forwards coins to the $\beta+1$ neighbors with the largest simulated layer. If more than $\beta+1$ neighbors have an $\infty$ layer, a node can forward the coins to any such $\beta+1$ neighbors. We show that with this forwarding rule, enough coins traverse to parts of $D(v)$ that have not received any coins before, and we keep learning more and more nodes of $D(v)$.
@@ 3.1 The AMPC model
In the AMPC model, we have an input of size $N$, which is processed by $P$ machines, each with space $S$. The unit of measure is a word of size $O(\log n)$ bits. The total space of all machines is $T=S\cdot P$ and we assume that $S=\Theta(N^{1-\Omega(1)})$. In addition, there is collection of distributed data stores (DDS) that we denote by $D_{0},D_{1},\dots$. We assume that all DDS provide a key-value store semantics: they store a collection of key-value pairs, such that given a key, a DDS returns the corresponding value. We require that both key and value consist of a constant number of words. The input data is stored in $D_{0}$ and uses a set of keys known to all machines (e.g. consecutive integers). The solution to a problem is stored in the last data store.
@@
The realism of the AMPC model is discussed in more depth in [${}^{+}$19]. The algorithms presented in this paper take a graph with $n$ nodes and $m$ edges as an input. In this case, the input size is $N=n+m$. We also assume that $S=n^{\delta}$, where $\delta\in(0,1)$ is some constant.
@@ 3.2 Arboricity
$2|E|=\sum_{v\in V}\deg(v)>\beta\cdot\frac{2\alpha|V|}{\beta}=2\alpha\cdot|V|\stackrel{{\scriptstyle*}}{{\geq}}2\cdot\frac{|E|}{|V|-1}\cdot|V|>2|E|\leavevmode\nobreak\ ,$
which is a contradiction. The inequality *** follows from Definition 3.1, as $|E|/(|V|-1)$ must be a lower bound for arboricity. ∎
@@ 3.3 $\beta$-Partition
Consider a graph $G=(V,E)$ and a function $\lambda\colon V(G)\xrightarrow{}\mathbb{N}\cup\{\infty\}$. For $i\in\mathbb{N}\cup\{\infty\}$ define $V_{i}=\{v\in V(G)\mid\lambda(v)=i\}$ and $V_{\geq i}=\bigcup_{j\geq i}V_{j}$. We say that $\lambda$ is a $\beta$-partition if $\deg_{V_{\geq i}}(v)\leq\beta$ for every node $v\in G$ with $\lambda(v)\neq\infty$ and $|V_{\infty}|=0$. If $|V_{\infty}|>0$, we call $\lambda$ a partial $\beta$-partition. We refer to $V_{i}$ as layer $i$ of the $\beta$-partition and $\lambda(v)$ as the layer of a node $v$. The size of a (partial) $\beta$-partition is the number of unique non-infinity layers.
@@
We extend the definition of $\beta$-partitions further to node-induced subgraphs, where its aprior unclear how to deal with neighbors not in the subgraph.
@@ 3.4 Induced $\beta$-partition
Barenboim and Elkin [BE10b] define an algorithm where we iteratively put all nodes with degree at most $(2+\varepsilon)\alpha$ in the top-most layer and delete them from the graph. This algorithm has the property that each node will have at most $(2+\varepsilon)\alpha$ neighbors in the same layer or above, and therefore, the output of this algorithm is an $(2+\varepsilon)\alpha$-partition. Using Facts 3.2 and 3.3, it is easy to see that size of the partition produced by this algorithm is at most $O(\log_{(2+\varepsilon)/2}n)$. In Definition 3.6, we generalize the algorithm of [BE10b] in two ways: first, by making the degree bound $\beta$ and second, by running it only on a subset $S$ of the nodes.
@@
    For $i=0,\dots,|S|$: Set $\sigma_{S,\beta}(v)\xleftarrow{}i$ for all nodes $v\in S$ with at most $\beta$ nodes $u\in N(v)$ for which $\sigma_{S,\beta}(u)=\infty$.
@@
In the following, an iteration refers to one round of Step 2 in Definition 3.6. Let $\sigma_{S}^{k}(v)$ denote value $\sigma_{S}(v)$ at the beginning of iteration $k$. For readability, we write $\sigma$ instead of $\sigma_{S,\beta}$. i. Observe that if there is an iteration where no node is assigned a $\sigma$ value, none of the following iterations assign $\sigma$ values. Hence, if value $|S|$ is assigned, $\sigma(v)\neq\infty$ holds for all nodes $v\in S$. Towards a contradiction, consider an iteration $i<|S|$ after which $\sigma(v)=\infty$ and there are at most $\beta$ nodes $w\in N(v)$ with $\sigma(w)=\infty$. In iteration $i+1$, it must be that $\sigma(v)\xleftarrow{}i+1\neq\infty$, a contradiction. ii. Consider $\sigma(v)=\sigma^{i}(v)=i\neq\infty$. By definition, it holds that $\sigma^{i}(w)=\infty$ for at most $\beta$ neighbors $w\in N(v)$. Hence, for said neighbors it will hold that $\sigma(w)=\infty$ or $\sigma(w)\in[i,|S|]$. In either case, $\sigma(w)\geq\sigma(v)$. iii. Consider $\sigma(v)=\sigma^{i}(v)=i\neq\infty$. Since $\deg(u)\geq\beta+1$, it must hold that $i>0$. Define $I_{i}$ to be the set of nodes $w\in N(v)$ for which $\sigma^{i}(w)=\infty$. Define $I_{i-1}$ similarly and observe that $I_{i}\subsetneq I_{i-1}$: since $\sigma(v)=i$, it must be that $|I_{i-1}|\geq\beta+1$ and $|I_{i}|\leq\beta$. Hence, for nodes $w\in I_{i-1}\setminus I_{i}$ it holds that $\sigma(w)=i-1$. Since $\sigma^{i}(w)=\infty$ for nodes $w\in I_{i}$, it must be that $\sigma(w)\geq i$. Hence, for $|I_{k-1}|\geq\beta+1$ nodes $w\in N(v)$ it holds that $\sigma(w)\geq i-1=\sigma(v)-1$. ∎
@@
For readability, we write $\sigma_{S}$ and $\sigma_{T}$ instead of $\sigma_{S,\beta}$ and $\sigma_{T,\beta}$, respectively. If $S=T$, then $\sigma_{S}$ equals $\sigma_{T}$ by definition. Hence, consider only $S\subset T$. The claim holds for all nodes $v$ for which $\sigma_{S}(v)=\infty$. Since $\sigma_{S}(v)=\infty$ holds for all nodes $v\in V\setminus S$ by definition, we only need to prove the claim for nodes $v\in S$ for which $\sigma_{S}(v)\neq\infty$. Consider the construction of the induced partitions $\sigma_{S}$ and $\sigma_{T}$ from Definition 3.6 side-by-side. Towards a contradiction, let $i$ be the first iteration (side-by-side) after which $\sigma_{S}(v)<\sigma_{T}(v)$ holds for a node $v\in S$. It must be that $\sigma_{S}(v)=i\neq\infty$ and $\sigma_{T}(v)=\infty$, which implies that $\sigma_{S}(w)=\infty$ for at most $\beta$ nodes $w\in N(v)$ and that $\sigma_{T}(v)=\infty$ for at least $\beta+1$ nodes $w\in N(v)$. Hence there must have existed one node $x$ for which $\sigma_{S}(x)<\sigma_{T}(x)$ after iteration $i-1$, a contradiction. ∎
@@ 3.5 Natural $\beta$-partition
Later, in our algorithmic usage of induced $\beta$-partitions, we will compute many of these partitions in parallel on various induced subgraphs $G[S]$, $S\subseteq V$ of the input graph $G$. The next lemma states that the level assignment of a node $v$ in $\sigma_{S,\beta}$ corresponds to its layer assignment in the natural $\beta$-partition if $G[S]$ contains the dependency graph of $v$.
@@
For readability, we write $\sigma_{S}$ and $\ell$ instead of $\sigma_{S,\beta}$ and $\ell_{\beta}$, respectively. Let us first prove the claim for $S=D(\ell,v)$ by induction on the values of $\ell(v)$, the base case being $\ell(w)=0$. The base case implies that $\deg(w)\leq\beta$, so by Definition 3.6 it holds that $\sigma_{S}(w)=0$. Assume that $\sigma_{S}(w)=\ell(w)$ for all nodes $w\in S$ with $\ell(w)<i$. Consider a node $u\in S$ with layer $\ell(u)=i$. Observe that all nodes $w\in N(u)$ with $\ell(w)<\ell(u)=i$ are in $S$ and $\sigma_{S}(w)=\ell(w)$ by induction. Since $\ell(u)=i$, by Lemma 3.7, there are at most $\beta$ nodes $w\in N(u)$ with $\ell(w)\geq\ell(u)$, implying that $\sigma_{S}(u)$ will be $i$. We have established that $\sigma_{D(\ell,v)}(w)=\ell(w)$ for every node $w\in D(\ell,v)$ if $S=D(\ell,v)$. If $S\supset D(\ell,v)$, the previous implies $\sigma_{S}(w)\leq\ell(w)$, because by Lemma 3.8 it holds that $\sigma_{S}(w)\leq\sigma_{D(\ell,v)}(w)$. On the other hand, by Lemma 3.13, for any subgraph $S\subseteq G$ it holds that $\sigma_{S}(w)\geq\ell(w)$ for every node $w\in S$. Hence, if $D(\ell,v)\subseteq S$ it holds that $\sigma_{S}(w)=\ell(w)$ for every node $w\in D(\ell,v)$. ∎

### v2

@@ 2 Sublinear LCA: High Level Challenges and Techniques
We design a deterministic LCA with a sublinear number of queries to find an acyclic orientation with a small out-degree $\beta=\Omega(\alpha)$ for a very large fraction of the nodes in a graph with arboricity $\alpha$. The informal statement is Lemma 1.1 and the formal one is Lemma 4.7.
@@ 2.1 Where should one query?
In order to steer the direction of the queries, consider the following coin dropping procedure. Give $n^{\delta}$ coins to node $v$, who naïvely (this approach will still not work) forwards the coins to all of its neighbors, giving each neighbor an equal share of the total number of $v$’s coins. Consider repeating this process recursively for every node holding coins, until coins cannot be further divided, which is in some sense a "volume-based" BFS. The forwarding of coins represents querying for nodes, and every node that has received a coin has been queried by $v$ and hence known to $v$. Optimally, all nodes of $D(v)$ receive a coin. Unfortunately, this approach fails for a large family of counter-examples: skewed dependency graphs (Figure 2b), comprising of a long path where every node has, e.g., $n^{\delta/10}$, neighbors in the lowest layer. In this case, $n^{\delta}$ coins will be already spent after only 10 steps of distributing coins, effectively learning only roughly $10\cdot n^{\delta/10}$ dependency graph nodes.
@@ 2.2 Our solution
Our solution uses the same coin dropping procedure, but we devise more sophisticated forwarding rules that determine how to distribute the coins. These rules utilize the topology of the graph induced by $v$’s knowledge aka the nodes $v$ has discovered so far. The forwarding rules are also updated frequently during the course of the algorithm to steer the future exploration of the graph and making very strong usage of the adaptivity of the queries.
Recall that every node in $D(v)$ has at most $\beta$ neighbors outside of $D(v)$. Hence, when coins are at a node $u\in D(v)$ and $u$ forwards coins to $\beta+1$ neighbors, we ensure that at least one neighbor from $D(v)$ receives coins, or more precisely at least a $1/(\beta+1)$-fraction of the coins goes to neighbors in $D(v)$. If we keep recursively applying this approach to all nodes that still have coins available, this roughly tells us, that after $k$ steps of forwarding, at least $x/(\beta+1)^{k}$ coins are at nodes that are contained in the dependency graph. Here $x$ is the number of coins that $v$ started with. However, it is not sufficient to choose $\beta+1$ neighbors arbitrarily, since it may be that we keep forwarding coins to nodes that we already know, i.e., nodes that have seen a coin before. In the end, it may be that all of these coins are just accumulated in a small part of $D(v)$ that we already know, or in other words, the process does not lead $v$ to querying new, undiscovered, parts of the dependency graph.
@@
Hence, we need to choose the correct neighbors to whom we forward the coins such that enough coins "flow" into the part of the dependency graph we have not yet learned, and not too many "leak out" outside of the dependency graph or "get stuck" in a part of the dependency graph that we already know. The rules for choosing the correct neighbors are chosen adaptively and frequently updated. Node $v$ keeps simulating an estimation of the $\beta$-partition (this estimation is called the induced $\beta$-partition, formalized in Definition 3.6) for the whole subgraph it has queried so far, which we call $S_{v}$ (which is only a small part of the whole graph). In the simulation, every node in $S_{v}$ is assigned a layer by $v$, or if no layer is assigned, we assume it is $\infty$. During an iteration, nodes forward coins to the $\beta+1$ neighbors with the largest simulated layer. If more than $\beta+1$ neighbors have an $\infty$ layer, a node can forward the coins to any such $\beta+1$ neighbors. We show that with this forwarding rule, enough coins traverse to parts of $D(v)$ that have not received any coins before, and we keep learning more and more nodes of $D(v)$.
@@ 3.1 The AMPC model
In the AMPC model, we have an input of size $N$, which is processed by $P$ machines, each with space $S$. The unit of measure is a word of size $O(\log n)$ bits. The total space of all machines is $T=S\cdot P$ and we assume that $S=\Theta(N^{1-\Omega(1)})$. In addition, there is a collection of distributed data stores (DDS) that we denote by $D_{0},D_{1},\dots$. We assume that all DDS provide a key-value store semantics: they store a collection of key-value pairs, such that given a key, a DDS returns the corresponding value. We require that both key and value consist of a constant number of words. The input data is stored in $D_{0}$ and uses a set of keys known to all machines (e.g. consecutive integers). The solution to a problem is stored in the last data store.
@@
The realism of the AMPC model is discussed in more depth in [BDE+19]. The algorithms presented in this paper take a graph with $n$ nodes and $m$ edges as an input. In this case, the input size is $N=n+m$. We also assume that $S=n^{\delta}$, where $\delta\in(0,1)$ is some constant.
@@ 3.2 Arboricity
$2|E|=\sum_{v\in V}\deg(v)>\beta\cdot\frac{2\alpha\cdot|V|}{\beta}=2\alpha\cdot|V|\stackrel{{\scriptstyle*}}{{\geq}}2\cdot\frac{|E|}{|V|-1}\cdot|V|>2|E|\leavevmode\nobreak\ ,$
which is a contradiction. The inequality * follows from Definition 3.1, as $|E|/(|V|-1)$ must be a lower bound for arboricity. ∎
@@ 3.3 $\beta$-Partition
Consider a graph $G=(V,E)$ and a function $\lambda\colon V(G)\xrightarrow{}\mathbb{N}\cup\{\infty\}$. For $i\in\mathbb{N}\cup\{\infty\}$ define $V_{i}=\{v\in V(G)\mid\lambda(v)=i\}$ and $V_{\geq i}=\bigcup_{j\geq i}V_{j}$.
We say that $\lambda$ is a $\beta$-partition if $\deg_{V_{\geq i}}(v)\leq\beta$ for every node $v\in G$ with $\lambda(v)\neq\infty$ and $|V_{\infty}|=0$. If $|V_{\infty}|>0$, we call $\lambda$ a partial $\beta$-partition. We refer to $V_{i}$ as layer $i$ of the $\beta$-partition and $\lambda(v)$ as the layer of a node $v$. The size of a (partial) $\beta$-partition is the number of unique non-infinity layers.
@@
We extend the definition of $\beta$-partitions further to node-induced subgraphs, where it is a priori unclear how to deal with neighbors not in the subgraph.
@@ 3.4 Induced $\beta$-partition
Barenboim and Elkin [BE10b] define an algorithm where we iteratively put all nodes with degree at most $(2+\varepsilon)\alpha$ in the top-most layer and delete them from the graph. This algorithm has the property that each node will have at most $(2+\varepsilon)\alpha$ neighbors in the same layer or above, and therefore, the output of this algorithm is a $(2+\varepsilon)\alpha$-partition. Using Facts 3.2 and 3.3, it is easy to see that size of the partition produced by this algorithm is at most $O(\log_{(2+\varepsilon)/2}n)$. In Definition 3.6, we generalize the algorithm of [BE10b] in two ways: first, by making the degree bound $\beta$ and second, by running it only on a subset $S$ of the nodes.
@@
    For $i=0,\dots,|S|$:
    Set $\sigma_{S,\beta}(v)\xleftarrow{}i$ for all nodes $v\in S$ with at most $\beta$ nodes $u\in N(v)$ for which $\sigma_{S,\beta}(u)=\infty$.
@@
In the following, an iteration refers to one round of Step 2 in Definition 3.6. Let $\sigma_{S}^{k}(v)$ denote value $\sigma_{S}(v)$ at the beginning of iteration $k$. For readability, we write $\sigma$ instead of $\sigma_{S,\beta}$.
i. Observe that if there is an iteration where no node is assigned a $\sigma$ value, none of the following iterations assign $\sigma$ values. Hence, if value $|S|$ is assigned, $\sigma(v)\neq\infty$ holds for all nodes $v\in S$. Towards a contradiction, consider an iteration $i<|S|$ after which $\sigma(v)=\infty$ and there are at most $\beta$ nodes $w\in N(v)$ with $\sigma(w)=\infty$. In iteration $i+1$, it must be that $\sigma(v)\xleftarrow{}i+1\neq\infty$, a contradiction.
ii. Consider $\sigma(v)=\sigma^{i}(v)=i\neq\infty$. By definition, it holds that $\sigma^{i}(w)=\infty$ for at most $\beta$ neighbors $w\in N(v)$. Hence, for said neighbors it will hold that $\sigma(w)=\infty$ or $\sigma(w)\in[i,|S|]$. In either case, $\sigma(w)\geq\sigma(v)$.
iii. Consider $\sigma(v)=\sigma^{i}(v)=i\neq\infty$. Since $\deg(u)\geq\beta+1$, it must hold that $i>0$. Define $I_{i}$ to be the set of nodes $w\in N(v)$ for which $\sigma^{i}(w)=\infty$. Define $I_{i-1}$ similarly and observe that $I_{i}\subsetneq I_{i-1}$: since $\sigma(v)=i$, it must be that $|I_{i-1}|\geq\beta+1$ and $|I_{i}|\leq\beta$. Hence, for nodes $w\in I_{i-1}\setminus I_{i}$ it holds that $\sigma(w)=i-1$. Since $\sigma^{i}(w)=\infty$ for nodes $w\in I_{i}$, it must be that $\sigma(w)\geq i$. Hence, for $|I_{k-1}|\geq\beta+1$ nodes $w\in N(v)$ it holds that $\sigma(w)\geq i-1=\sigma(v)-1$. ∎
@@
For readability, we write $\sigma_{S}$ and $\sigma_{T}$ instead of $\sigma_{S,\beta}$ and $\sigma_{T,\beta}$, respectively. If $S=T$, then $\sigma_{S}$ equals $\sigma_{T}$ by definition. Hence, consider only $S\subset T$. The claim holds for all nodes $v$ for which $\sigma_{S}(v)=\infty$. Since $\sigma_{S}(v)=\infty$ holds for all nodes $v\in V\setminus S$ by definition, we only need to prove the claim for nodes $v\in S$ for which $\sigma_{S}(v)\neq\infty$.
Consider the construction of the induced partitions $\sigma_{S}$ and $\sigma_{T}$ from Definition 3.6 side-by-side. Towards a contradiction, let $i$ be the first iteration (side-by-side) after which $\sigma_{S}(v)<\sigma_{T}(v)$ holds for a node $v\in S$. It must be that $\sigma_{S}(v)=i\neq\infty$ and $\sigma_{T}(v)=\infty$, which implies that $\sigma_{S}(w)=\infty$ for at most $\beta$ nodes $w\in N(v)$ and that $\sigma_{T}(v)=\infty$ for at least $\beta+1$ nodes $w\in N(v)$. Hence there must have existed one node $x$ for which $\sigma_{S}(x)<\sigma_{T}(x)$ after iteration $i-1$, a contradiction. ∎
@@ 3.5 Natural $\beta$-partition
Later, in our algorithmic usage of induced $\beta$-partitions, we will compute many of these partitions in parallel on various induced subgraphs $G[S]$, $S\subseteq V$ of the input graph $G$. The next lemma states that the layer assignment of a node $v$ in $\sigma_{S,\beta}$ corresponds to its layer assignment in the natural $\beta$-partition if $G[S]$ contains the dependency graph of $v$.
@@
For readability, we write $\sigma_{S}$ and $\ell$ instead of $\sigma_{S,\beta}$ and $\ell_{\beta}$, respectively.
Let us first prove the claim for $S=D(\ell,v)$ by induction on the values of $\ell(v)$, the base case being $\ell(w)=0$. The base case implies that $\deg(w)\leq\beta$, so by Definition 3.6 it holds that $\sigma_{S}(w)=0$. Assume that $\sigma_{S}(w)=\ell(w)$ for all nodes $w\in S$ with $\ell(w)<i$. Consider a node $u\in S$ with layer $\ell(u)=i$. Observe that all nodes $w\in N(u)$ with $\ell(w)<\ell(u)=i$ are in $S$ and $\sigma_{S}(w)=\ell(w)$ by induction. Since $\ell(u)=i$, by Lemma 3.7, there are at most $\beta$ nodes $w\in N(u)$ with $\ell(w)\geq\ell(u)$, implying that $\sigma_{S}(u)$ will be $i$.
We have established that $\sigma_{D(\ell,v)}(w)=\ell(w)$ for every node $w\in D(\ell,v)$ if $S=D(\ell,v)$. If $S\supset D(\ell,v)$, the previous implies $\sigma_{S}(w)\leq\ell(w)$, because by Lemma 3.8 it holds that $\sigma_{S}(w)\leq\sigma_{D(\ell,v)}(w)$. On the other hand, by Lemma 3.13, for any subgraph $S\subseteq G$ it holds that $\sigma_{S}(w)\geq\ell(w)$ for every node $w\in S$. Hence, if $D(\ell,v)\subseteq S$ it holds that $\sigma_{S}(w)=\ell(w)$ for every node $w\in D(\ell,v)$. ∎